OpenStudy (anonymous):

solve for x 1/4x-2=1/2x-3

OpenStudy (gw2011):

I assume you mean the following: (1/4)(x)-2=(1/2)(x)-3 (1/4)(x)-(1/2)(x)=-3+2=-1 Multiply both sides by 4 which eliminates the denominator x-2x=-4 -x=-4 x=4